OpenStudy (tmagloire1):

Find the x-coordinates where f '(x) = 0 for f(x) = 2x + sin(2x) in the interval [0, 2π]

OpenStudy (irishboy123):

go ahead and differentiate

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

did you find \[f'(x)\]?

OpenStudy (tmagloire1):

f'(x) =4cos^2(x)

OpenStudy (misty1212):

hmm no

OpenStudy (misty1212):

\[ f(x) = 2x + \sin(2x) \] right?

OpenStudy (tmagloire1):

Yep

OpenStudy (misty1212):

what is the derivative of \(2x\)?

OpenStudy (tmagloire1):

2

OpenStudy (misty1212):

ok good

OpenStudy (misty1212):

now how about the derivative of \(\sin(2x)\)?

OpenStudy (tmagloire1):

2cos2x

OpenStudy (misty1212):

ok

OpenStudy (misty1212):

so the derivative is not \(4\cos^2(x)\) is it?

OpenStudy (misty1212):

it is \[f'(x)=2+2\cos(2x)\]

OpenStudy (misty1212):

final job is to set \[2+2\cos(2x)=0\] and solve for \(x\)

OpenStudy (misty1212):

you good from there?

OpenStudy (tmagloire1):

Yes thank you! I was just getting the wrong derivative then my bad!

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (tmagloire1):

actually can someone help me find 2+2cos(2x)=0

zepdrix (zepdrix):

`subtract` 2cos(2x)=-2 `divide 2` cos(2x)=-1

zepdrix (zepdrix):

cosine of angle is -1 when the angle isssss...... pi, ya?

OpenStudy (tmagloire1):

Ohh ok thank you!

zepdrix (zepdrix):

careful though! the angle is (2x). that leads to 2x=pi still need to go further to solve for x.

OpenStudy (tmagloire1):

im still confused though. if it is 2x=pi than how do you get the x coordinates for the original problem @zepdrix

zepdrix (zepdrix):

divide... by 2 0_o

OpenStudy (tmagloire1):

oh.my.god.im retarded lolol. so x=pi/2

OpenStudy (tmagloire1):

Wow how could i not see that. gr8 moments

zepdrix (zepdrix):

Oh oh we should be careful though...\[\large\rm \cos(2x)=-1\qquad\to\qquad 2x=\pi+2k \pi,\qquad k\in \mathbb Z\]

zepdrix (zepdrix):

It's not just pi, we can also add any multiple of 2pi to that angle, ya? It's like spinning an extra time around the circle and landing in the same location.

zepdrix (zepdrix):

This will be important because we might get several solutions within this interval.

zepdrix (zepdrix):

Dividing by 2,\[\large\rm x=\frac{\pi}{2}+k \pi\]

OpenStudy (tmagloire1):

So pi/2 and 5pi/2

zepdrix (zepdrix):

we're using integers for k. So k=0 gives us pi/2, looks good! k=1 gives us ... ? k=2 gives us 5pi/2 <- this is larger than 2pi, out of our interval, so we can throw it away :)

OpenStudy (tmagloire1):

k=1 will give us 3pi/2

OpenStudy (tmagloire1):

@zepdrix

zepdrix (zepdrix):

Good! :) Here is a graph to help you see what is going on. https://www.desmos.com/calculator/xm7sfacgsm

zepdrix (zepdrix):

I graphed the function 2x+sin(2x) from 0 to 2pi. You can clearly see that at pi/2 and 3pi/2 we have critical points!

zepdrix (zepdrix):

That's what f'(x)=0 gives us, critical points. locations where the slope of the function is zero. graphically it looks like the top of a hill, or the bottom of a valley.

OpenStudy (tmagloire1):

Ohh okay so critical points is a another way of looking at it! Thank you so much!

zepdrix (zepdrix):

yay team \c:/